# Cluster algebras and quantum cohomology rings: A-type

### v1

@@ 5.5.1 Case one: the node $v$ has one adjacent node
This is just a special case of (5.48) when $m=k+1$. ∎
@@ 5.5.2 Case two: the node $v$ has two adjacent nodes
Assume the equation holds for some $a=k-p$. Then when $k-p=a+1$, we can apply Equation (5.48) to $c_{t}^{S}(S_{k+1}/S_{p})$ and get
@@
$=c_{t}^{S}(S_{k+2}/S_{p+1})*\left(c_{t}^{S}(S_{k+1}/S_{p+1})*c_{t}^{S}(S_{p+1}/S_{p})-(-1)^{N_{p+1}+N_{p}}q_{p+1}c_{t}^{S}(S_{k+1}/S_{p+2})\right).$ (5.51)
Again by Equation (5.48), the first term is
$c_{t}^{S}(S_{k+1}/S_{p+1})*\left(c^{S}_{t}(S_{k+2}/S_{p})+(-1)^{N_{p+1}+N_{p}}q_{p+1}c_{t}^{S}(S_{k+2}/S_{p+2})\right).$ (5.52)
@@
The second situation for a node admitting two valences is when there is a path passing through the node and no other arrows are adjacent to the node. There are two cases, the first is when $m=k+1,\,p=k-1$ and the second case is when $p=k-1,\,l=k-1$. Then the theorem 5.2 is reduced to the following Lemma.
@@
$c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{k+1}/S_{k})=c_{t}^{S}(S_{k+1}/S_{l})+(-1)^{N_{k-1}+N_{k}}q_{k}c_{t}^{S}(S_{k-1}/S_{l}),\,$ (5.55a)
$c_{t}^{S}(S_{k}/S_{k-1})*c_{t}^{S}(S_{m}/S_{k})=c_{t}^{S}(S_{m}/S_{k-1})+(-1)^{N_{k}+N_{k-1}}q_{k}c_{t}^{S}(S_{m}/S_{k+1})$ (5.55b)
The index $l$ in Equation (5.55a) is allowed to be zero.
@@
The Equation (5.55b) is exactly the Equation (5.48). We only have to prove the Equation (5.55a). We prove it by induction on the difference $k-l$. When $k-l=1$, it is exactly the equation in Lemma 5.13.
@@
$c_{t}^{S}(S_{l})*c_{t}^{S}(S_{k+1}/S_{l})*c_{t}^{S}(S_{k+2}/S_{k+1})=c_{t}^{S}(S_{l})*\left[c_{t}^{S}(S_{k+2}/S_{l})+(-1)^{N_{k}+N_{k+1}}q_{k+1}c_{t}^{S}(S_{k}/S_{l})\right]$ (5.56)
By utilizing (5.38), we expand the left hand side to
$c_{t}^{S}(S_{k+2})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k})+(-1)^{N_{l}+N_{l-1}}q_{l}c_{t}^{S}(S_{l-1})*\left[c_{t}^{S}(S_{k+1}/S_{l+1})*c_{t}^{S}(S_{k+2}/S_{k+1})\right].$ (5.57)
Again, by utilizing Equation (5.38), we expand the quantum product $c_{t}^{S}(S_{k+2}/S_{l})*c_{t}^{S}(S_{l})$ and $c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{l})$ and rewrite the right hand side as
$c_{t}^{S}(S_{k+2})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k})$
$+(-1)^{N_{l}+N_{l-1}}q_{l}c_{t}^{S}(S_{l-1})*\left[c_{t}^{S}(S_{k+2}/S_{l+1})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{l+1})\right].$ (5.58)
The terms inside brackets in Equations (5.57) and (5.5.2) are equal by the induction assumption. Hence the two sides of (5.56) are equal and the Lemma has been proved. ∎
@@
When a node has exactly three adjacent nodes, two of them are in a 3-cycle, and the remaining one is not in any $3$-cycle. The local behavior of the quiver diagram can be classified to the following three cases.
@@
$c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{m}/S_{l+1})=c_{t}^{S}(S_{m}/S_{l})*c_{t}^{S}(S_{k}/S_{l+1})+(-1)^{N_{k}+N_{l}}(\prod_{a=l+1}^{k}q_{a})c_{t}^{S}(S_{m}/S_{k+1})\,.$ (5.61)
In all above three equations, the index $l$ can be zero.
@@
When $k=p+1$, the equation is reduced to the Equation (5.55a) which has been proved to be true.
Assume that the equation holds for $k$. For $k+1$, utilizing (5.55a) we get
@@
$=c_{t}^{S}(S_{k+1}/S_{l})*\left(c_{t}^{S}(S_{k+2}/S_{k+1})*c_{t}^{S}(S_{k+1}/S_{p+1})-(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{p+1})\right)\,.$ (5.62)
Again by applaying (5.55a), we rewrite the first term as
$c_{t}^{S}(S_{k+1}/S_{p+1})*\left(c_{t}^{S}(S_{k+2}/S_{l})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{l})\right).$ (5.63)
Combine (5.5.3) and (5.63), and we get
@@
The proof is similar with that of (5.59). One can prove by induction on $m$. When $m=k+1$, the Equation (5.60) is reduced to (5.55a), and the Equation (5.61) is reduced to a special case of Equation (5.59). Assume that they hold for index smaller than or equal to $m$. For $m+1$, one needs to repeatedly use the Equation (5.55a) and prove the two equations. We leave the detail to readers. ∎
@@
We prove by induction on $m$. When $m=k+1$, the equation is reduced to the Equation (5.59). Assume that the Equation 5.65 holds for the index less than or equal to $m$. For $m+1$, we expand the $c_{t}^{S}(S_{m+1}/S_{p+1})$ by Equation (5.55a). Then we have
$c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{m+1}/S_{p+1})$
$=c_{t}^{S}(S_{k}/S_{l})\left(c_{t}^{S}(S_{m+1}/S_{m})*c_{t}^{S}(S_{m}/S_{p+1})-(-1)^{N_{m}+N_{m-1}}q_{m}c_{t}^{S}(S_{m-1}/S_{p+1})\right)$ (5.66)
By applying the induciton assumption for $m$, the first term is equal to
$c_{t}^{S}(S_{m+1}/S_{m})*\left(c_{t}^{S}(S_{m}/S_{l})*c_{t}^{S}(S_{k}/S_{p+1})+(-1)^{N_{p}+N_{k}}\prod_{a=p+1}^{k}q_{a}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})\right)$ (5.67)
The first term of (5.67) again by Equation (5.55a) can be rewritten as
$c_{t}^{S}(S_{k}/S_{p+1})*c_{t}^{S}(S_{m+1}/S_{l})+(-1)^{N_{m}+N_{m-1}}q_{m}c_{t}^{S}(S_{m-1}/S_{l})*c_{t}^{S}(S_{k}/S_{p+1}).$ (5.68)
The second term of (5.68) and the second term of the right hand side of the Equation (5.5.4), by the induction assumption on $m-1$, is equal to
$-(-1)^{N_{p}+N_{k}+N_{m}+N_{m+1}}(q_{m}\prod_{a=p+1}^{k}q_{a})c_{t}^{S}(S_{m-1}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l}).$ (5.69)
The second term of the formula (5.67), by using Equation (5.55a) is equal to
$(-1)^{N_{p}+N_{k}}(\prod_{a=p+1}^{k}q_{a})c_{t}^{S}(S_{p}/S_{l})*\left(c_{t}^{S}(S_{m+1}/S_{k+1})+(-1)^{N_{m}+N_{m+1}}q_{m}c_{t}^{S}(S_{m-1}/S_{k+1})\right)$ (5.70)
whose second term cancel with the formula in (5.69). Combine (5.5.4)-(5.70), we have proved the Equation (5.65). ∎
@@
The images $\{c_{t}^{S}(S_{k}/S_{l})\}$ of cluster variables for $A$-type quivers coincide with the conjectural images of cluster variables in the quantum cohomology rings for general quivers in Conjecture 3.8. That is to say the Chern polynomials of quotient bundles $c_{t}^{S}(S_{k}/S_{l})$ are pullbacks of Chern polynomials of the dual of tautological bundles over varieties of $QPs$ in $\Omega^{A}_{n}$.
@@
$c_{t}^{S}(S_{k}/S_{l})\big{|}_{P}=c_{t}^{S}(\tilde{S}_{v})\big{|}_{P^{\prime}}\,,$ (5.72)
where $P^{\prime}=(I_{u}^{\prime})_{u\in Q_{0}\backslash Q_{f}}$ is the image of $P$ in $(\mathcal{Z})^{S}$ by performing the rule in Lemma 4.7 along the quiver mutations $\mu_{v_{i}}\cdots\mu_{v_{1}}$. Denote the set of equivariant weights of of tautological bundles of flag variety $S_{i}|_{P}$ by $\mathcal{I}_{i}$, for $i=1,2,\ldots,n+1$, and the set of weights $\tilde{S}_{v}|_{P^{\prime}}$ by $\mathcal{I}_{v}^{\prime}$. Due to the rule in Lemma 4.7, $\mathcal{I}_{v}^{\prime}$ can be obtained from $\{\mathcal{I}_{i}\}_{i=1,\ldots,n+1}$ via a composition of disjoint unions and complements. Furthermore, the composition procedure is independent of the choice of integers $N_{1},\ldots,N_{n+1}$. For arbitrary $N_{1},\ldots,N_{n+1}$, we have $|\mathcal{I}_{i}|=N_{i}$, and $|\mathcal{I}^{\prime}_{v}|=N_{k}-N_{l}$, this implies $\mathcal{I}_{v}^{\prime}=\mathcal{I}_{k}\backslash\mathcal{I}_{l}$. So the (5.72) is proved.
@@
$c_{t}^{S}(\tilde{S}_{v})*c_{t}^{S}(\tilde{Q}_{v})=c_{t}^{S}(\tilde{S}_{v_{1}})*c_{t}^{S}(\tilde{S}_{v_{4}}^{*})+(-1)^{N_{k}-N_{l}}\tilde{q}_{v}c_{t}^{S}(\tilde{S}_{v_{2}}^{*})*c_{t}^{S}(\tilde{S}_{v_{3}})\,.$ (5.73)
where $Q_{v}$ is defined in Section 3.2. Let $\tilde{\beta}=(\tilde{\beta}_{u})\in H_{2}(\mathcal{Z},\mathbb{Z})$ and $\beta=(\beta_{i})\in H_{2}(Fl,\mathbb{Z})$ be the dual basis of first Chern classes of tautological bundles, as defined in (4.11). The $\tilde{q}_{v}$ in (5.73) is the Kähler variable tracking $\tilde{\beta}_{v}$. By (5.71), the equation (5.73) holds if we can prove that
$\phi_{*}(\tilde{\beta}_{v})=\beta_{p+1}+\cdots+\beta_{k},$ (5.74)
@@
$\int_{\tilde{\beta}_{v}}c_{1}(\tilde{S}_{w})=\int_{\beta_{p+1}+\cdots+\beta_{k}}\Phi^{-1}(c_{1}(\tilde{S}_{w}))$ (5.75)
For $w=v,v_{1},v_{2},v_{3},v_{4}$, the equation (5.75) is easily checked by (5.71) and (4.11). For other nodes, we use the following observation implied by Lemma 2.15. For each nodes $u,w\in\mathbf{Q}$, the subscripts $a,b,c,d$ of the associated integer $r_{w}=N_{a}-N_{b}$, $r_{u}=N_{c}-N_{d}$ satisfying the following condition
$\begin{cases}d\leq a,b\leq c,&\text{ if }w\prec u\\ a,b\notin(d,c)&\text{ otherwise }\,.\end{cases}$ (5.76)
By applying (5.76) for nodes $u=v_{1},v_{2},v_{3},v_{4}$, we can easily check the equation (5.75) by (5.71) and (4.11). ∎

### v2

@@ 5.5.1 Case one: the node $v$ has one adjacent node
This is just a special case of (64) when $m=k+1$. ∎
@@ 5.5.2 Case two: the node $v$ has two adjacent nodes
Assume the equation holds for some $a=k-p$. Then when $k-p=a+1$, we can apply Equation (64) to $c_{t}^{S}(S_{k+1}/S_{p})$ and get
@@
$=c_{t}^{S}(S_{k+2}/S_{p+1})*\left(c_{t}^{S}(S_{k+1}/S_{p+1})*c_{t}^{S}(S_{p+1}/S_{p})-(-1)^{N_{p+1}+N_{p}}q_{p+1}c_{t}^{S}(S_{k+1}/S_{p+2})\right).$
Again by Equation (64), the first term is
$c_{t}^{S}(S_{k+1}/S_{p+1})*\left(c^{S}_{t}(S_{k+2}/S_{p})+(-1)^{N_{p+1}+N_{p}}q_{p+1}c_{t}^{S}(S_{k+2}/S_{p+2})\right).$
@@
The second situation for a node admitting two valences is when there is a path passing through the node and no other arrows are adjacent to the node. There are two case. The first is when $m=k+1,\,p=k-1$ and the second case is when $p=k-1,\,l=k-1$. Then the theorem 5.2 is reduced to the following Lemma.
@@
(65a) $c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{k+1}/S_{k})=c_{t}^{S}(S_{k+1}/S_{l})+(-1)^{N_{k-1}+N_{k}}q_{k}c_{t}^{S}(S_{k-1}/S_{l}),\,$
(65b) $c_{t}^{S}(S_{k}/S_{k-1})*c_{t}^{S}(S_{m}/S_{k})=c_{t}^{S}(S_{m}/S_{k-1})+(-1)^{N_{k}+N_{k-1}}q_{k}c_{t}^{S}(S_{m}/S_{k+1})$
The index $l$ in Equation (65a) is allowed to be zero.
@@
The Equation (65b) is exactly the Equation (64). We only have to prove the Equation (65a). We prove it by induction on the difference $k-l$. When $k-l=1$, it is exactly the equation in Lemma 5.13.
@@
$c_{t}^{S}(S_{l})*c_{t}^{S}(S_{k+1}/S_{l})*c_{t}^{S}(S_{k+2}/S_{k+1})$
(66) $=c_{t}^{S}(S_{l})*\left[c_{t}^{S}(S_{k+2}/S_{l})+(-1)^{N_{k}+N_{k+1}}q_{k+1}c_{t}^{S}(S_{k}/S_{l})\right]$
By utilizing (57), we expand the left hand side to
(67) $c_{t}^{S}(S_{k+2})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k})+(-1)^{N_{l}+N_{l-1}}q_{l}c_{t}^{S}(S_{l-1})*\left[c_{t}^{S}(S_{k+1}/S_{l+1})*c_{t}^{S}(S_{k+2}/S_{k+1})\right].$
Again, by utilizing Equation (57), we expand the quantum product $c_{t}^{S}(S_{k+2}/S_{l})*c_{t}^{S}(S_{l})$ and $c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{l})$ and rewrite the right hand side as
$c_{t}^{S}(S_{k+2})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k})+(-1)^{N_{l}+N_{l-1}}q_{l}c_{t}^{S}(S_{l-1})*$
(68) $\left[c_{t}^{S}(S_{k+2}/S_{l+1})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{l+1})\right].$
The terms inside brackets in Equations (67) and (5.5.2) are equal by the induction assumption. Hence the two sides of (5.5.2) are equal and the Lemma has been proved. ∎
@@
When a node has exactly three adjacent nodes, two of them are in a 3-cycle, and the remaining one is not in any $3$-cycle. The local behavior of the quiver diagram can be classified into the following three cases.
@@
$c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{m}/S_{l+1})$
(71) $=c_{t}^{S}(S_{m}/S_{l})*c_{t}^{S}(S_{k}/S_{l+1})+(-1)^{N_{k}+N_{l}}(\prod_{a=l+1}^{k}q_{a})c_{t}^{S}(S_{m}/S_{k+1})\,.$
In all the above three equations, the index $l$ can be zero.
@@
When $k=p+1$, the equation is reduced to the Equation (65a) which has been proved to be true.
Assume that the equation holds for $k$. For $k+1$, utilizing (65a) we get
@@
(72) $=c_{t}^{S}(S_{k+1}/S_{l})*\left(c_{t}^{S}(S_{k+2}/S_{k+1})*c_{t}^{S}(S_{k+1}/S_{p+1})-(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{p+1})\right)\,.$
Again by applaying (65a), we rewrite the first term as
(73) $c_{t}^{S}(S_{k+1}/S_{p+1})*\left(c_{t}^{S}(S_{k+2}/S_{l})+(-1)^{N_{k+1}+N_{k}}q_{k+1}c_{t}^{S}(S_{k}/S_{l})\right).$
Combine (5.5.3) and (73), and we get
@@
The proof is similar with that of (5.16). One can prove by induction on $m$. When $m=k+1$, the Equation (5.16) is reduced to (65a), and the Equation (5.16) is reduced to a special case of Equation (5.16). Assume that they hold for index smaller than or equal to $m$. For $m+1$, one needs to repeatedly use the Equation (65a) and prove the two equations. We leave the detail to readers. ∎
@@
We prove by induction on $m$. When $m=k+1$, the equation is reduced to the Equation (5.16). Assume that the Equation 5.17 holds for the index less than or equal to $m$. For $m+1$, we expand the $c_{t}^{S}(S_{m+1}/S_{p+1})$ by Equation (65a). Then we have
$c_{t}^{S}(S_{k}/S_{l})*c_{t}^{S}(S_{m+1}/S_{p+1})=c_{t}^{S}(S_{k}/S_{l})*$
(76) $\left(c_{t}^{S}(S_{m+1}/S_{m})*c_{t}^{S}(S_{m}/S_{p+1})-(-1)^{N_{m}+N_{m-1}}q_{m}c_{t}^{S}(S_{m-1}/S_{p+1})\right)$
By applying the induction assumption for $m$, the first term is equal to
(77) $c_{t}^{S}(S_{m+1}/S_{m})*\left(c_{t}^{S}(S_{m}/S_{l})*c_{t}^{S}(S_{k}/S_{p+1})+(-1)^{N_{p}+N_{k}}\prod_{a=p+1}^{k}q_{a}c_{t}^{S}(S_{m}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l})\right)$
The first term of (77) again by Equation (65a) can be rewritten as
(78) $c_{t}^{S}(S_{k}/S_{p+1})*c_{t}^{S}(S_{m+1}/S_{l})+(-1)^{N_{m}+N_{m-1}}q_{m}c_{t}^{S}(S_{m-1}/S_{l})*c_{t}^{S}(S_{k}/S_{p+1}).$
The second term of (78) and the second term of the right hand side of the Equation (5.5.4), by the induction assumption on $m-1$, is equal to
(79) $-(-1)^{N_{p}+N_{k}+N_{m}+N_{m+1}}(q_{m}\prod_{a=p+1}^{k}q_{a})c_{t}^{S}(S_{m-1}/S_{k+1})*c_{t}^{S}(S_{p}/S_{l}).$
The second term of the formula (77), by using Equation (65a) is equal to
(80) $(-1)^{N_{p}+N_{k}}(\prod_{a=p+1}^{k}q_{a})c_{t}^{S}(S_{p}/S_{l})*\left(c_{t}^{S}(S_{m+1}/S_{k+1})+(-1)^{N_{m}+N_{m+1}}q_{m}c_{t}^{S}(S_{m-1}/S_{k+1})\right)$
whose second term cancels with the formula in (79). Combine (5.5.4)-(80), we have proved the Equation (5.17). ∎
@@
The images $\{c_{t}^{S}(S_{k}/S_{l})\}$ of cluster variables for $A$-type quivers coincide with the conjectural images of cluster variables in the quantum cohomology rings for general quivers in Conjecture 3.8. That is to say, the Chern polynomials of quotient bundles $c_{t}^{S}(S_{k}/S_{l})$ are pullbacks of Chern polynomials of the dual of tautological bundles over varieties of $QPs$ in $\Omega^{A}_{n}$.
@@
(82) $c_{t}^{S}(S_{k}/S_{l})\big{|}_{P}=c_{t}^{S}(\tilde{S}_{v})\big{|}_{P^{\prime}}\,,$
where $P^{\prime}=(I_{u}^{\prime})_{u\in Q_{0}\backslash Q_{f}}$ is the image of $P$ in $(\mathcal{Z})^{S}$ by performing the rule in Lemma 4.7 along the quiver mutations $\mu_{v_{i}}\cdots\mu_{v_{1}}$. Denote the set of equivariant weights of of tautological bundles of flag variety $S_{i}|_{P}$ by $\mathcal{I}_{i}$, for $i=1,2,\ldots,n+1$, and the set of weights $\tilde{S}_{v}|_{P^{\prime}}$ by $\mathcal{I}_{v}^{\prime}$. Due to the rule in Lemma 4.7, $\mathcal{I}_{v}^{\prime}$ can be obtained from $\{\mathcal{I}_{i}\}_{i=1,\ldots,n+1}$ via a composition of disjoint unions and complements. Furthermore, the composition procedure is independent of the choice of integers $N_{1},\ldots,N_{n+1}$. For arbitrary $N_{1},\ldots,N_{n+1}$, we have $|\mathcal{I}_{i}|=N_{i}$, and $|\mathcal{I}^{\prime}_{v}|=N_{k}-N_{l}$, this implies $\mathcal{I}_{v}^{\prime}=\mathcal{I}_{k}\backslash\mathcal{I}_{l}$. So the (82) is proved.
@@
(83) $c_{t}^{S}(\tilde{S}_{v})*c_{t}^{S}(\tilde{Q}_{v})=c_{t}^{S}(\tilde{S}_{v_{1}})*c_{t}^{S}(\tilde{S}_{v_{4}}^{*})+(-1)^{N_{k}-N_{l}}\tilde{q}_{v}c_{t}^{S}(\tilde{S}_{v_{2}}^{*})*c_{t}^{S}(\tilde{S}_{v_{3}})\,.$
where $Q_{v}$ is defined in Section 3.2. Let $\tilde{\beta}=(\tilde{\beta}_{u})\in H_{2}(\mathcal{Z},\mathbb{Z})$ and $\beta=(\beta_{i})\in H_{2}(Fl,\mathbb{Z})$ be the dual basis of first Chern classes of tautological bundles, as defined in (18). The $\tilde{q}_{v}$ in (83) is the Kähler variable tracking $\tilde{\beta}_{v}$. By (81), the equation (83) holds if we can prove that
$\phi_{*}(\tilde{\beta}_{v})=\beta_{p+1}+\cdots+\beta_{k},$
@@
(84) $\int_{\tilde{\beta}_{v}}c_{1}(\tilde{S}_{w})=\int_{\beta_{p+1}+\cdots+\beta_{k}}\Phi^{-1}(c_{1}(\tilde{S}_{w}))$
For $w=v,v_{1},v_{2},v_{3},v_{4}$, the equation (84) is easily checked by (81) and (18). For other nodes, we use the following observation implied by Lemma 2.15. For each nodes $u,w\in\mathbf{Q}$, the subscripts $a,b,c,d$ of the associated integer $r_{w}=N_{a}-N_{b}$, $r_{u}=N_{c}-N_{d}$ satisfying the following condition
(85) $\begin{cases}d\leq a,b\leq c,&\text{ if }w\prec u\\ a,b\notin(d,c)&\text{ otherwise }\,.\end{cases}$
By applying (85) for nodes $u=v_{1},v_{2},v_{3},v_{4}$, we can easily check the equation (84) by (81) and (18). ∎
[…]
Weiqiang He, Department of Mathematics, Sun Yat-sen University, Xingang West Road No.135, Haizhu District, Guangzhou, China.
Email: hewq@mail2.sysu.edu.cn.
Yingchun Zhang, Room 221, School of Mathematical Sciences, Shanghai Jiao Tong University, Shanghai, China.
Email: yingchun.zhang@sjtu.edu.cn